OpenStudy (anonymous):

what is the solution of the system of equations? -3x-4y-3z=-7 2x-6y+2z=3 5x-2y+5z=9 and could i please get a step by step solution so i can really teach myself how it was done? thanks!! the choices are (5,-2,7), (-5,2,7) (5,2,-7) and no solution

OpenStudy (anonymous):

please help im really confused!

OpenStudy (anonymous):

You want to reduce the number of variables by adding the equations together. Start by multiply eq. 1 by 2/3 and add it to eq.2

OpenStudy (anonymous):

why by 2/3?

hero (hero):

The easiest way to solve this is to try plugging in the points.

OpenStudy (anonymous):

is it like (x,y,z) ?

hero (hero):

Also, as a possible hint, for each equation the co-efficient of the x variable is the same as the coefficient for the z variable.

hero (hero):

This is not the best problem to do for teaching yourself how to solve systems of equations in three variables. Doing this one will only confuse you.

hero (hero):

You need to try systems that actually have solutions.

hero (hero):

But I will provide you with a step by step solution since that is what you ask for.

OpenStudy (anonymous):

okay im going to really still try to teach my self still because im still struggling and i need to know how to do these so itll really help thank you!

hero (hero):

Pair the 1st and 3rd equations together: -3x-4y-3z=-7 5x-2y+5z=9 Pair the 2nd and 3rd equations together. 2x-6y+2z=3 5x-2y+5z=9

OpenStudy (anonymous):

ohh, okay thank you so much! im gonna take note of this.

hero (hero):

For the 1st and 3rd equations, multiply 5x - 2y + 5z = 9 by -2 -3x - 4y - 3z = -7 -10x + 4y -10z = -18 Combine both equations together to get: -13x - 13z = -25 For the 2nd and 3rd equations, multiply 5x -2y + 5z = 9 by -3 2x-6y+2z=3 -15x+ 6y-15z=-27 Combine both together to get -13x -13z = -24 So after all that you end up with -13x - 13z = -25 -13x - 13z = -24 This is not possible.